# Precovery Observations of 3I/ATLAS from TESS Suggests Possible Distant Activity

### v1

@@ Abstract
3I/ATLAS is the third macroscopic interstellar object detected traversing the Solar System. Since its initial discovery on UT 01 July 2025, hundreds of hours on a range of observational facilities have been dedicated to measure the physical properties of this object. These observations have provided astrometry to refine the orbital solution, photometry to measure the color, a rotation period and secular light curve, and spectroscopy to characterize the composition of the coma. Here, we report precovery photometry of 3I/ATLAS as observed with NASA’s Transiting Exoplanet Survey Satellite (TESS). 3I/ATLAS was observed nearly continuously by TESS from UT 07 May 2025 to 02 June 2025. We use the shift-stack method to create deep stack images to recover the object. These composite images reveal that 3I/ATLAS has an average TESS magnitude of and an absolute visual magnitude of , consistent with magnitudes reported in July 2025, suggesting that 3I/ATLAS may have been active out at au. Additionally, we extract a day light curve and find no statistically significant evidence of a nucleus rotation period. Nevertheless, the data presented here are some of the earliest precovery images of 3I/ATLAS and may be used in conjunction with future observations to constrain the properties of our third interstellar interloper.
@@ 2.1 FFI Selection and Background Removal
We downloaded all FFIs for Camera 2 CCD 3 and Camera 1 CCD 2 using the bulk download scripts provided by the Mikulski Archive for Space Telescopes (MAST)222 https://archive.stsci.edu/missions-and-data/tess. 3I/ATLAS was observed on pixels on Camera 2 CCD 3 and pixels on Camera 1 CCD 2. We highlight example images of these two regions in Figure 1. It is important to note that columns are serial register columns, columns are virtual columns, and rows are buffer, smear, and virtual rows; none of these regions can be used for science.
The TESS FFIs are contaminated by both zodiacal light and scattered earthshine contamination. In particular, the scattered earthshine manifests as strong ramps at the beginning and end of each TESS orbit ( days). While the background typically varies smoothly, there have been instances of more sporadic and variable behavior (Figure 2). To identify highly contaminated FFIs, we calculated the median value in each FFI . The FFIs containing highly non-smooth variability are entirely removed from our analysis (Figure 2). This resulted in the removal of 9% of the available FFIs from Camera 2 CCD and 5% of the available FFIs from Camera 1 CCD 2. In total, after these data-quality cuts, we analyzed 9837 FFIs (353 GB). We provide more specific details pertaining to each Camera/CCD configuration in Table 1.
Once we removed heavily contaminated FFIs, we fit and remove the background from the remaining FFIs. Working with the entire FFI is unnecessary. Instead, we create a “postcard” cutout (A. D. Feinstein et al., 2019) of the entire region where 3I/ATLAS is visible in the FFI (orange boxes in the upper two panels and full bottom two panels in Figure 1). Previous search and recovery analyses of known solar system bodies with TESS have fit pixel variability using an nth-order polynomial (e.g. C. Kiss et al., 2025). We choose to model each pixel within the postcard with a data-driven Savistky-Golay filter with a window length of 307 exposures and a 2nd-order polynomial. We opt for this more flexible model as it allows us to accurately model the steep earthshine ramps, without removing any sharp features which may be astrophysical in origin. We explored using smaller and larger window-sizes. We found that smaller window sizes removed astrophysical signals, while larger windows were unable to capture the variability of the orbital ramps.
@@
| [au] | 5.82 - 5.33 | 5.19 - 4.60 |
| 4399 | 5438 |
@@ 2.2 Deep Image Creation
After the background models are removed, we employ a shift-stack algorithm to create a TESS deepstack image. We choose not to combine images across detectors, as they may have different systematics. Our algorithm works as follows: (I) Query the location of 3I/ATLAS from the JPL Horizons Small Bodies Database for each FFI; (II) Use the world coordinate system information to convert the queried (RA, Dec) to the (x,y) pixel location of 3I/ATLAS; (III) Create a pixel cutout around where 3I/ATLAS is expected to be; (IV) Repeat this procedure for all integrations; (V) Sum all of the images. Additionally, due to the apparent location with respect to the galactic plane, we create a data quality flag to track particularly crowded fields. FFIs are flagged when at least of the pixels in the cutout are from the median background value. Our methodology works well for objects with known orbital parameters, but is significantly more computationally inefficient when trying to search for new objects. However, it has been employed with some success for example by W. C. Fraser et al. (2024) — albeit using more sensitive data. A coarse summation presented here is sufficient for TESS due to its large pixel scale.
@@
We present the deepstack image of 896 Sphinx in Figure 3. We find that we are able to recover 896 Sphinx at . Additionally, we extract a light curve using a -pixel mask centered on 896 Sphinx. We present the extracted light curve in Figure 3. We find signatures of variability ranging from hours, which is consistent with the known rotation period (T. Polakis, 2018). We note that our algorithm takes minutes to recover this object.
We present the two deepstack images of 3I/ATLAS as observed by TESS in Figure 4. We stack all available FFIs for Camera 2 CCD 3. We mask the neighboring one pixel around 3I/ATLAS and calculate the median and standard deviation of the background. We find that we are able to detect 3I/ATLAS at above the median background. We do not stack all available FFIs for Camera 1 CCD 2, as the field becomes increasingly more crowded as 3I/ATLAS’ apparent motion places it closer to the galactic plane. We use our data quality flag and only use frames that are not significantly contaminated with nearby stars. We stacked 2518 exposures to produce this image (Figure 4, right). We repeat the same significance test and find that we detect 3I/ATLAS at with respect to the median background. Both images represent statistically significant detects of 3I/ATLAS in the TESS FFIs.
@@ 3.1 Photometric Light Curve
We use a 1-pixel aperture to extract a TESS light curve for 3I/ATLAS. Additionally, we create a running-median light curve with a bin width of , which roughly corresponds to 36 minutes in each bin. We convert our light curve from e- s-1 to TESS magnitude, using the following relationship:
In Equation 1, is the number of counts per second. This conversion is provided in the TESS Instrumental Handbook (R. Vanderspek et al., 2018). It is important to note that the zero-point has an uncertainty of 0.05 mag. 3I/ATLAS has a in the Camera 2 CCD 3 image and a in the Camera 1 CCD 2 image. From geometric effects alone, and ignoring the small () phase angle change, we would expect the flux from 3I to increase by a factor of 1.5. Converting this back into magnitudes we expected 3I/ATLAS to have a measured magnitude of 19.9 in the Camera 1 CCD 2 images. Within our uncertainty on the magnitude, our observations are statistically inconsistent with an asteroid-like reflectance model that ignores dust production.
These two magnitudes are averaged over the two observing windows in Table 1, which span a range in heliocentric distance of 6.4 to 5.4 au, respectively. In that same time period the distance between 3I/ATLAS and the TESS spacecraft decreased by 0.9 au. The rapid change in both these properties can be used to constrain any changes to the overall activity of 3I/ATLAS between the two time periods by investigating if the measured 0.1 mag decrease is consistent with the expected change from both the decreased heliocentric and TESS distances.
@@
We convert our light curves into by multiplying the normalized light curve by the average magnitude and analyze our median light curve with a Lomb-Scargle periodogram (N. R. Lomb, 1976; J. D. Scargle, 1982) to search for evidence of periodicity. We do this for each light curve independently. We search for a rotation period between hours, which is mostly consistent with of other kilometer scale small bodies in the solar system (B. D. Warner et al., 2009). Both of the light curves and affiliated periodograms are presented in Figure 5.
@@
There are several similarities between the Lomb-Scargle periodograms. First, we find that hours are dominated by noise and/or unremoved systematics. Second, we see that there is no agreement between periodograms at hours. Third, we see a peak at hours in both periodograms. Although the peaks in the periodograms are offset by 1.3 hours, the similar signal from both detectors could be promising evidence of a rotation period. If hours, this would be discrepant from the current reported period of hours (R. de la Fuente Marcos et al., 2025).
@@
To further assess this potential rotational signature, we run the same Lomb-Scargle test for a nearby background pixel in the FFI. Because the background can rapidly vary across the TESS detector, we select an example background pixel within -pixels of 3I/ATLAS. The results of this test are presented in Figure 6. We find that the background pixel periodograms show a similar peak at hours. The background pixel peaks are offset from the 3I/ATLAS peaks by 1.3 and 1.6 hours in each image, respectively. Therefore, these TESS light curves do no definitively capture a from the nucleus. This may be due to ongoing outgassing even at large heliocentric distances or due to the sensitivity of TESS, making any possible magnitude changes from the baseline difficult to definitively measure.
@@ 3.2 Secular Light Curve
In Equation 3, is the comet-Sun distance, is the comet-Earth distance, and is an activity index. has a typical range from , where assumes the body is inactive and assumes the object has strong activity (E. Everhart, 1967). Given that 3I/ATLAS is at a distance of 6.4 au in these images, we assume .
@@ 4 Discussion
We present two deepstack images and a 20-day light curve of 3I/ATLAS as observed with NASA’s Transiting Exoplanet Survey Satellite. These data were obtained from 07 May to 03 June 2025, predating the initial discovery by nearly two months (L. Denneau et al., 2025). From these data we calculate that 3I/ATLAS had a , which is consistent with more recently obtained observations from the TRAPPIST-North and TRAPPIST-South telescopes (D. Z. Seligman et al., 2025). Moreover, we identify a consistent and periodic signal between both images at hours. However, we find no statistically significant evidence of a rotation period from the nucleus when comparing the light curve of 3I/ATLAS to a neighboring background pixel. The lack of observed rotation may be due to the presence of a coma, uncertainties in the size of the nucleus if there is no coma, or decreased sensitivity of TESS to faint objects.
@@
There is no evidence for significant changes to the absolute magnitude of 3I/ATLAS during these observations, at least within the sensitivity limits of the data. It is worth noting that the object traversed nearly 1 au during the course of these observations. This provides tentative evidence that 3I/ATLAS was active during the fortuitous TESS observations prior to its discovery on UT 01 July 2025. However, the 3 error technically permits the possibility that 3I/ATLAS was fainter by 1 magnitude during these observations. This is due to the uncertainties in the conversion from magnitudes to of 0.3 (R. Vanderspek et al., 2018; T. L. Farnham et al., 2021). Therefore, we conclude that these observations are consistent with the hypothesis that 3I/ATLAS was weakly active prior to its discovery. Future work could include deriving a new relationship between and other bandpasses for ISOs.
@@
This distant activity of 3I/ATLAS — if confirmed — could be indicative of mass loss driven by a mechanism other than the sublimation of H2O ice. One possibility is that 3I/ATLAS is enriched in hypervolatiles such as CO or CO2. 2I/Borisov doisplayed a higher production of CO than H2O when these two quantities were measured contemporaneously (D. Bodewits et al., 2020; M. A. Cordiner et al., 2020; Z. Xing et al., 2020), although no direct measurement of CO2 has been successfully obtained in an interstellar object to date.333Production rates of CO2 have been inferred from the ration of O I emission features in the visible, A. McKay et al. 2024 It is also possible that this activity is driven by a more exotic hypervolatile such as H2 or N2. The nongravitational acceleration of 1I/‘Oumuamua energetically would have required the sublimation of an ice more volatile than H2O (Z. Sekanina, 2019; D. Seligman & G. Laughlin, 2020). It has therefore been proposed that this acceleration was driven by outgassing of N2 (A. P. Jackson & S. J. Desch, 2021), CO (D. Z. Seligman et al., 2021), or radiolytically produced H2 (J. B. Bergner & D. Z. Seligman, 2023). It is possible that distant activity of 3I/ATLAS is driven by one of these, but determining which neutral is responsible requires spectroscopic confirmation or long-term lightcurve modeling (E. Bufanda et al., 2023).
@@
The relative flatness of the light curves could be interpreted to indicate very slow changes to the activity rate, which is likely due to the rapidly changing thermal environment of the object. Continued searches for precovery images and extensive monitoring of 3I/ATLAS prior to perihelion will be essential for characterizing this trend. Moreover, spectroscopic follow-up observations could also confirm the existence of hypervolatiles in 3I/ATLAS. Of particular interest in the precovery images would be any stellar occultations or appulses that may exist in the dataset (J. L. Ortiz et al., 2023; C. L. Pereira et al., 2025), especially from multi-filter observatories that could constrain coma densities and composition.

### v2

@@ Abstract
3I/ATLAS is the third macroscopic interstellar object detected traversing the Solar System. Since its initial discovery on UT 01 July 2025, hundreds of hours on a range of observational facilities have been dedicated to measure the physical properties of this object. These observations have provided astrometry to refine the orbital solution, photometry to measure the color, a rotation period and secular light curve, and spectroscopy to characterize the composition of the coma. Here, we report precovery photometry of 3I/ATLAS as observed with NASA’s Transiting Exoplanet Survey Satellite (TESS). 3I/ATLAS was observed nearly continuously by TESS from UT 07 May 2025 to 02 June 2025. We use the shift-stack method to create deep stack images to recover the object. These composite images reveal that 3I/ATLAS has an average TESS magnitude of and an absolute visual magnitude of , the latter being consistent with magnitudes reported in July 2025. When coupled with recent HST images deriving a nucleus size of R2.8 km (H15.4), our measurements suggest that 3I/ATLAS may have been active out at au. Additionally, we extract a day light curve and find no statistically significant evidence of a nucleus rotation period. Nevertheless, the data presented here are some of the earliest precovery images of 3I/ATLAS and may be used in conjunction with future observations to constrain the properties of our third interstellar interloper.
@@ 2.1 FFI Selection and Background Removal
We downloaded all calibrated FFIs for Camera 2 CCD 3 and Camera 1 CCD 2 using the bulk download scripts provided by the Mikulski Archive for Space Telescopes (MAST)222 https://archive.stsci.edu/missions-and-data/tess. 3I/ATLAS was observed on pixels on Camera 2 CCD 3 and pixels on Camera 1 CCD 2. We highlight example images of these two regions in Figure 1. It is important to note that columns are serial register columns, columns are virtual columns, and rows are buffer, smear, and virtual rows; none of these regions can be used for science.
The TESS FFIs are contaminated by both zodiacal light and scattered earthshine contamination. In particular, the scattered earthshine manifests as strong ramps at the beginning and end of each TESS orbit ( days). While the background typically varies smoothly, there have been instances of more sporadic and variable behavior (Figure 2). To identify highly contaminated FFIs, we calculated the median value in each FFI . The FFIs containing highly non-smooth variability, as identified by-eye, are entirely removed from our analysis (Figure 2). This resulted in the removal of 9% of the available FFIs from Camera 2 CCD and 5% of the available FFIs from Camera 1 CCD 2. In total, after these data-quality cuts, we analyzed 9837 FFIs (353 GB). We provide more specific details pertaining to each Camera/CCD configuration in Table 1.
We fit and subtract the background from all remaining FFIs after the heavily contaminated ones were removed. Working with the entire FFI is unnecessary. Instead, we create a “postcard” cutout (A. D. Feinstein et al., 2019) of the entire region where 3I/ATLAS is visible in the FFI (orange boxes in the upper two panels and full bottom two panels in Figure 1). Previous search and recovery analyses of known solar system bodies with TESS have fit pixel variability using an nth-order polynomial (e.g. C. Kiss et al., 2025). We experimented with two methods to remove background contamination from these observations. First, we fit a 2nd-order polynomial to the background, similar to C. Kiss et al. (2025). We fit each continuous section of the light curve with these models, i.e. we chunk the observations based on gaps in the TESS data. This leaves a total of five different segments that are fit (two segments for Camera 2 CCD 3; three segments for Camera 1 CCD 2). This experiment demonstrated that a 2nd-order polynomial fits the baseline well, but fails to fit the orbital ramps (Figure 10).
Second, we fit the background with a data-driven Savitsky-Golay filter with a pre-defined window length and a 2nd-order polynomial. We repeat the same process of chunking the light curve based on gaps in the observations. We test several window lengths ranging from 31 - 901 pixels. The residuals of several example model fits are presented in Figure 10. This experiment revealed that smaller window sizes removed astrophysical signals, while larger windows did not capture the variability of the orbital ramps. Therefore, we settled on using an average window length of 307 pixels.
@@
| [au] | 6.36 - 6.02 | 5.92 - 5.48 |
| 4399 | 5438 | |
2.2 Removing Crowded Frames
Due to the crowdedness in some of the images, we create a filter to remove images which may have significant stellar contamination. The final deepstack images in this work only include frames that are not significantly contaminated. The contamination criteria is defined as the following: Because we are only interested in crowding near the location of 3I/ATLAS, we only consider sources within an window around the central location. We use astropy.stats.sigma_clip to determine how many pixels are outliers compared to the mean value. We assume the default settings for astropy.stats.sigma_clip, other than defining the value. We mask any frame for which more than half of the pixels within this region are flagged as outliers. Examples of non-crowded versus crowded frames are shown in Figure 3.
@@ 2.3 Deep Image Creation
After the background models are removed, we employ a shift-stack algorithm to create a TESS deepstack image. We choose not to combine images across detectors, as they may have different systematics. Our algorithm works as follows: (I) Query the location of 3I/ATLAS from the JPL Horizons Small Bodies Database for each FFI; (II) Use the world coordinate system information to convert the queried (RA, Dec) to the (x,y) pixel location of 3I/ATLAS. We convert the (x, y) location from float values to integers using np.round, in order to accurately determine the nearest pixel; (III) Create a pixel cutout centered on the integer (x, y) coordinates where 3I/ATLAS is expected to be; (IV) Repeat this procedure for all integrations; (V) Sum all of the images. Additionally, due to the apparent location with respect to the galactic plane, we create a data quality flag to track particularly crowded fields. The shift-stack technique implemented in this paper works well for objects with known orbital parameters, but may be more computationally inefficient when completing a blind-search for new objects. However, it has been employed with some success for example by W. C. Fraser et al. (2024) — albeit using more sensitive data. A coarse summation presented here is sufficient for TESS due to its large pixel scale.
2.4 Recovery of Known Minor Planet 896 Sphinx
@@
We present the deepstack image of 896 Sphinx in Figure 4. We find that we are able to recover 896 Sphinx at . Additionally, we extract a light curve using a -pixel mask centered on 896 Sphinx. We present the extracted normalized light curve and associated errors in Figure 4. We run this light curve through a Lomb-Scargle periodogram (N. R. Lomb, 1976; J. D. Scargle, 1982) to recover the known rotation period of the object. We find that periods hours are dominated by instrumental noise. We ignore these periods. We find the peak of max power for hours. We fit this peak with a Gaussian function and optimize our fit using a non-linear least squares approach. We find that the peak occurs at hours. Our error is adopted from the width of the best-fit Gaussian. Our recovered period is within of the period presented in T. Polakis (2018). The discrepancy between our rotation period and the archival rotation period may be due to the lack of sensitivity of TESS for faint objects.
2.5 Recovery of 3I/ATLAS
Figure 5 shows the background-removed deepstack images and associated errors of 3I/ATLAS as observed by TESS. These deepstacks are limited to cutouts which are not overcrowded with bright sources, as defined in Section 2.2. The median, mean, and summed images are shown for both detectors. The detection threshold for 3I/ATLAS is calculated by propagating errors from the calibrated FFIs. Although the median deepstack images are implemented for the remainder of this paper, it should be noted that 3I/ATLAS is visible by-eye in all three types of images in Figure 5. The central pixel in Camera 2 CCD 3 has a median counts value of e- s-1. The background scatter is calculated as the median and standard deviation of the singular pixels surrounding the central pixel exercise. The background has an averaged count value of e- s-1. Therefore, we are able to recover 3I/ATLAS at in this summed deepstack image. The same exercise is repeated for Camera 1 CCD 2. The central pixel in this image has a summed counts value of e- s-1. The background has an averaged count value of e- s-1. Therefore, we are able to recover 3I/ATLAS at in this summed deepstack image. It is not surprising that we are unable to recover 3I/ATLAS at the same high fidelity as the previous image. This is because the field becomes increasingly more crowded as 3I/ATLAS’ apparent motion places it closer to the galactic plane.
@@ 3 Results
3.1 Magnitude Calculation
We use a pixel aperture on the median deepstacked image to calculate the magnitude of 3I/ATLAS. We convert the calibrated flux in e- s-1 to TESS magnitude using the following relationship:
In Equation 1, is the number of counts per second. This conversion is provided in the TESS Instrumental Handbook (R. Vanderspek et al., 2018). We adopt the error on the zeropoint of mag. We note that while the zeropoint may change slightly between sectors, cameras, and/or CCDs, it is standard practice to not recalculate the value for each observation combination (e.g. T. L. Farnham et al., 2019; M. M. Fausnaugh et al., 2023; J. E. Andrews et al., 2025). We calculate using a aperture on the median deepstack images presented in Figure 5. For Camera 2 CCD 3, we derive e- s-1. For Camera 1 CCD 2, we derive e- s-1.
We find that 3I/ATLAS has a in the Camera 2 CCD 3 image and a in the Camera 1 CCD 2 image. Our calculated is in agreement with J. Martinez-Palomera et al. (2025). Examining the origins of slight offsets between our extracted values is beyond the scope of this work. These two magnitudes are averaged over the two observing windows in Table 1, which span a range in heliocentric distance of 6.4 to 5.4 au, respectively. In that same time period the distance between 3I/ATLAS and the TESS spacecraft decreased by 0.9 au. The rapid change in both these properties can be used to constrain any changes to the overall activity of 3I/ATLAS between both times.
From geometric effects alone, and ignoring the small () phase angle change, we would expect the flux from 3I to increase by a factor of 1.5. Converting this back into magnitudes we expected 3I/ATLAS to have a measured magnitude of 20.5 in the Camera 1 CCD 2 images, rather than the observed of 19.28, indicating a factor of 5 increase in the flux, rather than 1.5. Within our uncertainty on the TESS magnitude, our observations are statistically inconsistent with an asteroid-like reflectance model that ignores dust production.
@@ 3.2 Photometric Light Curve
We extract a light curve using the same -pixel aperture and propagate the errors from the calibrated FFIs. We normalize the light curve. Additionally, we bin our light curve to 36-minute bins. We analyze the binned light curve with a Lomb-Scargle periodogram to search for evidence of periodicity. We do this for each light curve independently. We search for a rotation period between hours, which is mostly consistent with of other kilometer scale small bodies in the solar system (B. D. Warner et al., 2009). Both of the light curves and affiliated periodograms are presented in Figure 6.
@@
There are several similarities between the Lomb-Scargle periodograms. First, we find that hours are dominated by noise and/or unremoved systematics. Second, we see that there is no agreement between periodograms at hours. Third, we see a some strong periodicity between 10-30 hours. This is more apparent in the light curve from Camera 1 CCD 2, when 3I/ATLAS 0.4 au closer. In particular, we find a peak in the periodogram close to the recently recovered period of 16.16-16.79 hours presented in R. de la Fuente Marcos et al. (2025); T. Santana-Ros et al. (2025).
@@
To further assess this potential rotational signature, we run the same Lomb-Scargle test for a nearby background pixels in the FFI. Because the background can rapidly vary across the TESS detector, we select an example background pixel within -pixels of 3I/ATLAS. The results of this test are presented in Figure 7. We find that the background pixel periodogram for Camera 2 CCD 3 shows a similar peak at hours, but we do not see a similarly strong peak around this time for Camera 1 CCD 2. However, there are stronger periodic peaks in this periodogram, so it is challenging to assess whether or not this 16 hour period is physically motivated by the TESS observations alone.
We complete this exercise for all of the pixels within a region around 3I/ATLAS. We pay particular attention to pixels that lead and trail the source on sky, as they should reveal systematics produced by nearby stars. The leading and trailing pixels are defined as those diagonal to the center from the bottom left to the upper right (the direction of motion of 3I/ATLAS). The light curve extracted from each pixel is binned at a 36 minute resolution. A Lomb-Scargle periodogram is performed on each resulting light curve across the same frequency range (Figure 8). There are some similarities between the leading/trailing pixels that do not align with the periodic signals observed from 3I/ATLAS. In particular, the trailing/leading pixels display a similar periodic signal at 29 hours in the data from Camera 2 CCD 3. This signal is offset from a peak in the 3I/ATLAS observations by hours. There are similar offsets between the trailing/leading and 3I/ATLAS pixels for Camera 1 CCD 2.
Given the similarities in noise properties of the periodograms, it is challenging to draw any confident conclusions about 3I/ATLAS. Therefore, we conclude that these TESS light curves do no definitively capture a from the nucleus. This may be due to ongoing outgassing even at large heliocentric distances or due to the sensitivity of TESS, making any possible magnitude changes from the baseline difficult to definitively measure.
@@ 3.3 Secular Light Curve
In Equation 3, is the comet-Sun distance, is the comet-TESS distance, and is an activity index. has a typical range from , where assumes the body is inactive and assumes the object has strong activity (E. Everhart, 1967). Given that 3I/ATLAS is at a distance of 6.4 au in these images, we assume .
@@ 4 Discussion
We present two deepstack images and a 20-day light curve of 3I/ATLAS as observed with NASA’s Transiting Exoplanet Survey Satellite. These data were obtained from 07 May to 03 June 2025, predating the initial discovery by nearly two months (L. Denneau et al., 2025). From our two epochs of these data we calculate that 3I/ATLAS had a and , the latter of which is consistent with more recently obtained observations from the TRAPPIST-North and TRAPPIST-South telescopes (D. Z. Seligman et al., 2025). Moreover, we identify a periodic signal at hours in the light curve from Camera 1 CCD 2. However, we find no statistically significant evidence of a rotation period from the nucleus when comparing the light curve of 3I/ATLAS to a neighboring background pixel. The lack of observed rotation may be due to the presence of a coma, uncertainties in the size of the nucleus if there is no coma, or decreased sensitivity of TESS to faint objects.
@@
The derived from these data indicate that 3I/ATLAS brightened more than can be expected from the decrease in helio- and TESS-centric distances between the two periods. It is worth stating that the object traversed nearly 1 au during the course of these observations. We note that the absolute magnitude of 3I/ATLAS in our first epoch,=13.9, is consistent with activity from a small (R2.8 km, 15.4) nucleus derived from HST WFC3 imaging (D. Jewitt et al., 2025). This provides tentative evidence that 3I/ATLAS was active during the fortuitous TESS observations prior to its discovery on UT 01 July 2025. However, the 3 error technically permits the possibility that 3I/ATLAS was fainter by 1 magnitude during these observations. This is due to the uncertainties in the conversion from magnitudes to of 0.3 (R. Vanderspek et al., 2018; T. L. Farnham et al., 2021). Therefore, we conclude that these observations are consistent with the hypothesis that 3I/ATLAS was weakly active prior to its discovery. Future work could include deriving a new relationship between and other bandpasses for ISOs.
@@
This distant activity of 3I/ATLAS — if confirmed — could be indicative of mass loss driven by a mechanism other than the sublimation of H2O ice. One possibility is that 3I/ATLAS is enriched in hypervolatiles such as CO or CO2. 2I/Borisov displayed a higher production of CO than H2O when these two quantities were measured contemporaneously (D. Bodewits et al., 2020; M. A. Cordiner et al., 2020; Z. Xing et al., 2020), although no direct measurement of CO2 has been successfully obtained in an interstellar object to date.333Production rates of CO2 have been inferred from the ration of O I emission features in the visible, A. McKay et al. 2024 It is also possible that this activity is driven by a more exotic hypervolatile such as H2 or N2. The nongravitational acceleration of 1I/‘Oumuamua energetically would have required the sublimation of an ice more volatile than H2O (Z. Sekanina, 2019; D. Seligman & G. Laughlin, 2020). It has therefore been proposed that this acceleration was driven by outgassing of N2 (A. P. Jackson & S. J. Desch, 2021), CO (D. Z. Seligman et al., 2021), or radiolytically produced H2 (J. B. Bergner & D. Z. Seligman, 2023). It is possible that distant activity of 3I/ATLAS is driven by one of these, but determining which neutral is responsible requires spectroscopic confirmation or long-term lightcurve modeling (E. Bufanda et al., 2023).
@@
The change in absolute magnitude between our two deep stacks indicates a change in activity, likely the result of a rapidly warming thermal environment. Continued searches for precovery images and extensive monitoring of 3I/ATLAS prior to perihelion will be essential for characterizing this trend. In particular, other precovery images which may include stellar occultations of appulses (J. L. Ortiz et al., 2023; C. L. Pereira et al., 2025), especially from multi-filter observations, could constrain the coma density and composition. Moreover, spectroscopic follow-up observations could also confirm the existence of hypervolatiles in 3I/ATLAS.